### v1

@@ 3.4 Learned Complementary Selection
$\mathbf{P}_{\textrm{fuse}}=w\mathbf{P}_{\textrm{mono}}+(1-w)\mathbf{P}_{\textrm{mv}},$ (2)
where $0\leq w\leq 1$ is the output by the MLP, $\mathbf{P}_{\textrm{mono}}$ and $\mathbf{P}_{\textrm{mv}}$ denote the probability distributions from a single and multi view, respectively. The expectation of the fused probability distribution $\mathbf{P}_{\textrm{fuse}}$ then provides the final depth prediction.
@@ 3.5 Sequential Localization
$p(\mathbf{s}_{t}|\mathbf{o}_{t})=e^{-\|\hat{\mathbf{r}}-\mathbf{r}_{\mathbf{s}_{t}}\|_{1}},$ (3)
where $\mathbf{r}_{\mathbf{s_{t}}}$ is the floorplan ray at pose $\mathbf{s}_{t}$ and $\hat{\mathbf{r}}$ is the interpolated ray from the floorplan depth prediction. We use the relative pose between frames, i.e., ego-motion as the transition model
$\mathbf{s}_{t+1}=\mathbf{s}_{t}\oplus\mathbf{t}_{t}+\mathbf{\omega}_{t},$ (4)
@@
$p(\mathbf{s}_{t+1}|\mathbf{o}_{t},\mathbf{t}_{t})=\frac{1}{Z}\sum_{\mathbf{s}_{t}}{p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{t}_{t})p(\mathbf{s}_{t}|\mathbf{o}_{t})},$ (6)
where $Z$ is a normalization factor.
@@
$\delta\mathbf{t}=\mathbf{R}_{s_{\phi}}^{-1}[s_{x},s_{y}]^{\top}-[t_{x},t_{y}]^{\top}$ (8)
with $\mathbf{R}_{s_{\phi}}\in\mathbb{R}^{2\times 2}$ being the rotation matrix with angle $s_{\phi}$. The rotational filter $\mathbf{r}$ is
@@
Figure 5: Transition as grouped convolution. (a) Illustration of the translational filters (from left to right, top to bottom the filters for 0, 10 to 350°normal-$\degree$ ) and the rotational filters derived from a sample ego-motion. (b) The probability volume is divided into $O$ groups, where $O$ is the number of orientations. Each group is convolved with its respective translational filter and stacked back together. After circular padding along the orientation axis, the volume is convolved with the rotational filter to finish the transition step.
@@ 4.3 Training Scheme
Table 1: Comparison between our observation model and the baselines.
@@
Gibson(g) R@ 0.1m 0.5m 1m 1m30° PF-net 1.0 1.9 5.6 1.9 LASER 0.7 7.0 11.8 9.5 $\text{Ours}_{s}$ 4.3 26.7 33.7 32.3 $\text{Ours}_{m}$ 9.3 27.0 31.0 29.2 $\text{Ours}_{t}$ 10.5 34.3 39.6 38.0 $\text{Ours}_{f}$ 12.2 39.4 44.5 43.2
@@ 5.1 Observation Model
Single Frame. We evaluate single frame localization accuracy on Gibson(f) and Structured3D. As shown in Tab. 2, the proposed monocular network, $\text{Ours}_{s}$ significantly outperforms both baselines on Gibson(f) seeing almost 200% improvement across all metrics. Also on Structured3D our method surpasses the state of the art by a large margin. When taking the orientation into account, the recall does not drop much (35.1% at 1m 30 deg compared to 36.6% at 1m on Gibson(f) and 21.3% to 22.4% on Structured3D). This underlines the accurate orientation estimation of our method. We notice here the performance of LASER on Structured3D does not align with that reported in [28], we suspect this is due to slightly difference in the dataset (perspective Structured3D compared to their perspective images cropped from panoramic Structured3D) and the random roll pitch angles this dataset contains.
@@ 6 Real-world Experiment
Since no real-world indoor dataset with both sequential observations and floorplan exist that allows training and testing, we show the potential of our pipeline in real-world scenario by customizing LaMAR [35]. LaMAR is a real world dataset containing three scenes. We select the trajectories in HGE indoor scene containing trajectories within a single floor, and split it into training and testing set. We use our single frame observation model with the proposed histogram filter to localize. The entire floor has an area of 80m×$\times$120m, and the data includes challenging observations as shown in Fig. 11. We use the data within 75m×$\times$81m and localize within the floorplan of the same size. Our system localizes and tracks the camera pose from the second step and closely follows it afterwards. Despite the large scene scale, our histogram filter is still efficient enough to localize at 3 hz.

### v2

@@ 3.4 Learned Complementary Selection
$\mathbf{P}_{\textrm{fuse}}=w\text{Upsample}(\mathbf{P}_{\textrm{mono}})+(1-w)\mathbf{P}_{\textrm{mv}},$ (2)
where $0\leq w\leq 1$ is the output by the MLP, $\mathbf{P}_{\textrm{mono}}$ and $\mathbf{P}_{\textrm{mv}}$ denote the probability distributions from a single and multi view, respectively. The upsampling ensures the validity of the addition. The expectation of the fused probability distribution $\mathbf{P}_{\textrm{fuse}}$ then provides the final depth prediction.
@@ 3.5 Sequential Localization
$p(\mathbf{s}_{t}|\mathbf{o}_{t})=e^{-\frac{1}{\lambda}\|\hat{\mathbf{r}}-\mathbf{r}_{\mathbf{s}_{t}}\|_{1}},$ (3)
where $\mathbf{r}_{\mathbf{s_{t}}}$ is the floorplan ray at pose $\mathbf{s}_{t}$, $\hat{\mathbf{r}}$ is the interpolated ray from the floorplan depth prediction and $\lambda$ a constant factor. We use the relative pose between frames, i.e., ego-motion as the transition model
$\mathbf{s}_{t+1}=\mathbf{t}_{t}\oplus\mathbf{s}_{t}+\mathbf{\omega}_{t},$ (4)
@@
$p(\mathbf{s}_{t+1}|\mathbf{o}_{t},\mathbf{t}_{t})=\frac{1}{Z}\sum_{\mathbf{s}_{t}\in\mathcal{S}}{p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{t}_{t})p(\mathbf{s}_{t}|\mathbf{o}_{t})},$ (6)
where $Z$ is a normalization factor and $\mathcal{S}$ the poses space.
@@
$\delta\mathbf{t}=\mathbf{R}_{\phi}^{-1}[x,y]^{\top}-[t_{x},t_{y}]^{\top}$ (8)
with $\mathbf{R}_{\phi}\in\mathbb{R}^{2\times 2}$ being the rotation matrix with angle ${\phi}$. The rotational filter $\mathbf{r}$ is
@@
Figure 5: Transition as grouped convolution. (a) Illustration of the translational filters (from left to right, top to bottom the filters for 0, 10 to 350°$\degree$ ) and the rotational filters derived from a sample ego-motion. (b) The probability volume is divided into $O$ groups, where $O$ is the number of orientations. Each group is convolved with its respective translational filter and stacked back together. After circular padding along the orientation axis, the volume is convolved with the rotational filter to finish the transition step.
@@ 4.3 Training Scheme
Gibson(g)
R@ 0.1m 0.5m 1m 1m30°
PF-net 1.0 1.9 5.6 1.9
LASER 0.7 7.0 11.8 9.5
$\text{Ours}_{s}$ 4.3 26.7 33.7 32.3
$\text{Ours}_{m}$ 9.3 27.0 31.0 29.2
$\text{Ours}_{t}$ 10.5 34.3 39.6 38.0
$\text{Ours}_{f}$ 12.2 39.4 44.5 43.2
@@ 5.1 Observation Model
Single Frame. We evaluate single frame localization accuracy on Gibson(f) and Structured3D. As shown in Tab. 2, the proposed monocular network, $\text{Ours}_{s}$ significantly outperforms both baselines on Gibson(f) seeing almost 200% improvement across all metrics. Also on Structured3D our method surpasses the state of the art by a large margin. When taking the orientation into account, the recall does not drop much (35.1% at 1m 30 deg compared to 36.6% at 1m on Gibson(f) and 21.3% to 22.4% on Structured3D). This underlines the accurate orientation estimation of our method. We notice here the performance of LASER on Structured3D does not align with that reported in [28], we suspect this is due to the difference in the dataset (perspective Structured3D compared to their perspective images cropped from panoramic Structured3D) and the random roll pitch angles this dataset contains.
@@ 6 Real-world Experiment
Since no real-world indoor dataset with both sequential observations and floorplan exist that allows training and testing, we show the potential of our pipeline in real-world scenario by customizing LaMAR [35]. LaMAR is a real world dataset containing three scenes. We select the trajectories in HGE indoor scene containing trajectories within a single floor, and split it into train and test set. We use our single frame observation model with the proposed histogram filter to localize. The entire floor has an area of 80m×$\times$120m, and the data includes challenging observations as shown in Fig. 11. We use the data within 75m×$\times$81m and localize within the floorplan of the same size. Our system localizes and tracks the camera pose from the second step and closely follows it afterwards. Despite the large scene scale, our histogram filter is still efficient enough to localize at 3 hz.